OpenStudy (bossimbacon):

Which ordered pairs in the form (x, y) are solutions to the equation 4x – 5y = 24? Choose all answers that are correct. A. (−9, −12) B. (1, −4) C. (4, 8) D. (6, 0) @midhun.madhu1987 @gorv

OpenStudy (midhun.madhu1987):

Just plug in each values of x and y in the equation and check whether it satisfies...

OpenStudy (bossimbacon):

@jaxiblu

OpenStudy (anonymous):

A. 4(-9) - 5(-12) = -36 - (-60) = 24 B. 4(1) - 5(-4) = 4 - (-20) = 24 C. 4(4) - 5(8) = 16 - 40 = -24 NOT 24 D. 4(6) - 5(0) = 24 - 0 = 24 So you'd choose A, B, and D.